OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x^2 - 5x + 6, find the function (f - g)(x). Use the ^ key for the exponent. Enter your answer as the following example: -2x^2+3x+10.

jhonyy9 (jhonyy9):

so what mean f-g(x) so just subtract the equation of(x) from equation of g(x) so this will be 5x-6 -(x^2 -5x +6) = 5x-6-x^2 +5x -6 = -x^2 +10x -12 hope this is understandably sure right easy good luck bye